# Complexity-aware fine-tuning

### v1

@@ Abstract
General-purpose Large Language Models (LLMs) are frequently fine-tuned through supervised fine-tuning (SFT) to enhance performance in specific domains. Better results can be achieved by distilling the chain-of-thought of a larger model at the cost of numerous expensive calls and a much greater amount of data. We propose a novel blueprint for efficient fine-tuning that uses reasoning only for complex data identified by entropy. Specifically, across two small open models ($\approx 3B$) we split the training data into complexity categories by a single token answer entropy (ROC AUC $0.73$), fine-tune large language models (LLMs) via SFT and distillation, and show that our pipeline significantly outperforms the standard SFT approach ($0.55$ vs $0.43$ average accuracy) and provides comparable with distillation performance while using $62\%$ less data ($0.55$ average accuracy for both). We publish our code [^1] https://github.com/LabARSS/complexity-aware-fine-tuning and data [^2] https://github.com/LabARSS/complexity-aware-fine-tuning?tab=readme-ov-file#data to facilitate further research in this direction.
@@
[…]
@@ 1 Introduction
LLMs demonstrate impressive performance across a vast variety of domains. At the same time, they show subpar results on niche tasks and in niche domains compared to specialized models. Properly tuned smaller models beat large open models in mathematics Yang et al. (2024b), medicine Wu et al. (2025), chemistry Yu et al. (2024), and other fields. While in some areas, the difference is negligible or at least not as important, in others, such as medicine, it becomes crucial due to the high cost of error.
A standard approach to get a domain-specific model is to fine-tune a general-purpose base LLM on carefully selected data Parthasarathy et al. (2024). While there is a lot of focus on perfecting the training mechanics, fewer attempts have been made to enhance the performance by pursuing meaningful data separation and ordering. The default approach is fine-tuning a model on the entire set without particular ordering. There have been experiments with curriculum-based learning Kim and Lee (2024); Shi et al. (2025), but they struggle to split the dataset by difficulty in a meaningful and automated fashion.
@@
Figure 1: Complexity-aware fine-tuning scheme: for easy and medium complex questions we apply direct SFT, while for hard questions we include reasoning results during SFT.
@@
To address this gap, we propose a fully automated pipeline, presented in Figure 1. It consists of two steps: (1) a split of the data for fine-tuning by the complexity of the task via a response entropy t get easy, medium and hard questions, and (2) training a model via supervised fine-tuning (SFT) followed by reasoning-promoting techniques (such as distillation Li et al. (2023); Hsieh et al. (2023) or, hypothetically, Group Relative Policy Optimization Zhihong Shao (2024); Face (2025)) for easy and hard data entries accordingly. We confirm the effectiveness of the framework by fine-tuning two open models: Qwen2.5-3B Yang et al. (2024a) and Phi-4-Mini Microsoft et al. (2025), on the multiple choice question answering dataset MMLU-Pro Wang et al. (2024).
@@
For both steps of the pipeline, we provide sensitivity studies to the design choices made. In step one, we consider different options for complexity estimation based on a model’s confidence in its answer. Our study considers a wide range of methods, including model-as-judge (MASJ), entropy-based aggregation, and calculation methods. We collect the generation results accompanied by token distribution at each step and publish them as an ensemble of open datasets for further research. In step two, we highlight that we need additional reasoning-based distillation only for hard questions. The specific contributions are the following:
  • An unsupervised method to split a multiple-choice question answering dataset by complexity based on the token-wise entropy of the response with ROC AUC $0.73$.
@@
  • An automated training pipeline to fine-tune LLMs via SFT and distillation for easy and hard questions accordingly. We show that our procedure significantly outperforms the standard SFT approach with accuracies $0.50$ and $0.60$ over two LLMs for MMLU-Pro c.t. $0.40$ and $0.46$ for baselines. It provides comparable performance with the distillation performance while using $62\%$ less data (accuracies $0.50$ and $0.60$ over two LLMs for MMLU-Pro c.t. $0.49$ and $0.62$ for baselines).
  • An analysis of complexity estimation rooted in MASJ-based and entropy-based estimates, augmented with reasoning results and various aggregation techniques. Non-aggregated entropy-based complexity yields the best results, with ROC AUC of $0.73$.
@@ 2 Related works
Fadeeva et al. (2023) provide an in-depth comparison of multiple black-box and white-box methods of complexity estimation. They present promising results for white-box methods rooted in sequence probability and entropy aggregation. Sychev et al. (2025) focus specifically on entropy-based aggregations augmented with model-as-judge (MASJ) categorization by a reasoning score. They confirm that LLM’s token-level entropy of the output is a good predictor of question difficulty, especially in knowledge-based domains. They also introduce MASJ reasoning score to estimate the question complexity. However, the authors use these metrics only to analyze model behavior. They do not integrate it into a practical data aggregation or fine-tuning workflow.
At the moment, it remains a mystery if we can combine novel complexity estimation techniques based on entropy aggregates with adaptive learning methods. It is also unclear how we can use the insights for the reasoning estimates to dynamically change our learning approach. This paper aims to address these gaps.
@@ 3.1 Training pipeline
We propose the complexity-aware fine-tuning pipeline (Figure 1) with the following major stages: complexity estimation, data aggregation, fine-tuning.
@@
To aggregate the data into groups by complexity (easy, medium, hard), we evenly divide the dataset into three parts, ordering the entries by entropy of the response. A group with the lowest entropy values is categorized as easy and with the highest values as hard, as Figure 2 depicts.
Figure 2: Data aggregation
@@
For easy and medium groups, we suggest using SFT Howard and Ruder (2018); Raffel et al. (2023) as an established practice to enhance the model performance. It involves fine-tuning a pretrained LM on labeled examples using standard supervised objectives (like cross-entropy). Here, we prompt (see Table 5) the model with the question and options. Then, we apply the cross-entropy loss to the question answers used as labels.
As to the hard group, we hypothesize that as hard questions require multiple logical steps, the model can not effectively learn with standard SFT and answer the question directly. Instead, we propose to elicit a chain-of-thought and allow the model to incrementally build the answer step-by-step as suggested by Wei et al. (2023).
In this work, we apply the distillation technique, which involves training a smaller model on the chain-of-thought of a larger LLM. It is a well-known practice supported by Hsieh et al. (2023). To create the distillation training samples, we prompt a large LLM to answer the multiple-choice question and produce a chain-of-thought in the process. Next, the whole response is attached to the dataset and used to train the smaller model. Hypothetically, other reasoning-eliciting techniques might show good performance at this stage as well.
@@ 3.2 Complexity estimation approaches
To find the most suitable metric for the training pipeline, we analyze the performance of the following techniques: MASJ reasoning score, MASJ education level, a single token answer entropy, a chain-of-thought answer entropy, a chain-of-thought aggregated response entropy.
Used prompts are available in Appendix A.1.
@@ 3.2.2 MASJ education level
3.2.3 Single token answer entropy
@@
where $p_{i}$ is the probability of a single token and $n$ is the vocabulary size.
@@
3.2.4 Chain-of-thought answer entropy
@@
3.2.5 Chain-of-thought aggregated response entropy
Building upon the single-token entropy approach, we investigate more sophisticated methods (in line with Fadeeva et al. (2023)) for complexity estimation by analyzing the entire chain-of-thought (CoT) response. While the answer token entropy provides a localized measure of uncertainty, aggregating entropy across the complete reasoning process potentially offers a more comprehensive complexity assessment.
@@ 4.1 Experimental setup
We conduct all experiments on the multiple choice question answering dataset MMLU-Pro Wang et al. (2024), widely adopted by the community as one of the main performance benchmarks. It spans 14 domains, offering a broad selection of questions of varying complexities. Each question has approximately 10 options, with a single correct one, which eliminates ambiguity in evaluation.
@@
Regarding the models  [^3] All models and datasets are published under permissive licenses that allow them to be used for research purposes, we utilize a variety of open model sizes for data collection and aggregation to analyze how the trend changes with model size. At the same time, we focus on smaller models for fine-tuning to make our results reproducible.
We apply our overall pipeline to two models: Qwen2.5-3B and Phi-4-Mini. For them, we measure single token and chain-of-thought entropy, collect other response metadata, fine-tune the models, and evaluate the overall pipeline.
Auxiliary models are used to extend our analysis and allow advanced reasoning: a single token response entropy - Mistral 24B and Phi-4, reasoning scores and education levels via MASJ with Mistral 123B Mistral (2024), chain-of-thought distillation - DeepSeek-V3-0324 DeepSeek-AI (2024).
We also conduct an additional experiment to collect metadata for a reasoning model using Qwen3-8B.
4.2 Fine-tuning
4.2.1 Complexity-aware pipeline
(a) Accuracy for fine-tuning pipelines after 10 epochs for Qwen 3B
Based on ROC AUC results and positive performance of SFT for medium and easy questions (see 4.2.2), we take single token entropy as our complexity metric for the pipeline described in Section 3.1.
@@ 4.2.1 Complexity-aware pipeline
As our baseline, we train the model via SFT without the data split for 10 epochs.
@@
Table  1 and Figures  3(a)  3(b) show the results. We see that the proposed training scheme results in a significant improvement over the baselines and an alternative training scheme that uses distillation for only easy and medium questions. Qwen 3B achieves an accuracy of 0.50 compared to 0.47 and 0.40 for alternative and baseline, respectively, while Phi-4-mini gets to 0.60 compared to 0.58 and 0.46. At the same time, we see that the pipeline provides comparable performance to the distillation with significantly less data (in number of tokens processed). Qwen 3B achieves an accuracy of 0.51 compared to 0.49 saving 59% data, while Phi-4-mini gets to 0.60 compared to 0.62 saving 64% data. We can also detect an uptrend in the pipeline training while the full distillation approach exhibits the signs of a plateau.
@@
SFT 0.40 / 3k 0.46 / 3k
Distillation 0.49 / 1,957k 0.62 / 1,482k
Alternative 0.47 / 1,158k 0.58 / 951k
Pipeline (ours) 0.51 / 801k 0.60 / 534k
@@
4.2.2 Sensitivity study
@@
The same data splitting procedure as described in 4.2.1 is applied for MASJ reasoning score and single-token entropy as a complexity metrics.
Figures 4(d) 4(c) show the results of SFT for each group split by the single-token entropy. For Phi-4-mini, we see that medium and easy questions outperform hard ones for the first 10 epochs. Shortly after, performance starts to decline for all groups. For Qwen 3B, we do not see a significant difference between the groups. Moreover, the performance plateaus after five epochs.
Figures 4(b) and 4(a) show the results of SFT for each group split by the MASJ reasoning score. We do not see a strong difference in performance between the groups. In combination with questionable ROC AUC scores, provided in Table 8, it makes MASJ reasoning score a less favorable metric for further experiments.
@@
4.3 Complexity estimation evaluation

### v2

@@ Abstract
General-purpose Large Language Models (LLMs) are frequently fine-tuned through supervised fine-tuning (SFT) to enhance performance in specific domains. Better results can be achieved by distilling the chain-of-thought of a larger model at the cost of numerous expensive calls and a much greater amount of data. We propose a novel blueprint for efficient fine-tuning that uses reasoning only for complex data identified by entropy. Specifically, across two small open models ($\approx 3B$) we split the training data into complexity categories by a single token answer entropy (ROC AUC $0.73$), fine-tune large language models (LLMs) via SFT and distillation, and show that our pipeline significantly outperforms the standard SFT approach ($0.58$ vs $0.45$ average accuracy) and outperforms the distillation approach ($0.58$ vs $0.56$ average accuracy) while using $81\%$ less data. We publish our code [^1] https://github.com/LabARSS/complexity-aware-fine-tuning/tree/eacl2026 and data [^2] https://github.com/LabARSS/complexity-aware-fine-tuning/tree/eacl2026?tab=readme-ov-file#data to facilitate further research in this direction.
@@
Andrey Goncharov1, Daniil Vyazhev1, Petr Sychev1, Edvard Khalafyan, Alexey Zaytsev1, 1Applied AI Institute
@@ 1 Introduction
General-purpose LLMs excel across diverse tasks, but deploying them is often impractical under constraints on compute, latency or cost. These realities motivate compact, domain-adapted models that can deliver competitive or superior performance. Growing evidence shows that carefully tuned smaller models can match or outperform larger open models in mathematics Yang et al. (2024b), medicine Wu et al. (2025), chemistry Yu et al. (2024). Thus, smaller domain-specific LLMs appear a compelling choice, especially under resource constraints.
A standard way for training is domain-adaptive fine-tuning: starting from a general-purpose LLM and training it on in-domain data Parthasarathy et al. (2024) with optional distillation from a larger model. Recent work has mainly concentrated on optimizing training mechanics (objectives, schedulers, and regularizers), whereas comparatively little attention has been paid to data curation. Curriculum-style approaches have also been explored Kim and Lee (2024); Shi et al. (2025) with increasing complexity of data during training, but with limited gains in final accuracy and efficiency, suggesting that more principled strategies are needed.
@@
Figure 1: Complexity-aware fine-tuning scheme for a student LLM: we identify complexity of questions via uncertainty estimation of a model (Step 1), then for questions of regular complexity we apply direct SFT (Step 2), while for hard questions we include reasoning from a teacher LLM (Step 3) and complete SFT using reasoning-enriched hard data (Step 4).
@@
To address this gap, we propose a fully automated pipeline, presented in Figure 1, that focus on both accuracy and efficiency of the fine-tuning procedure. It consists of two steps: (1) a split of the data for fine-tuning by the complexity of the task to regular and hard questions and (2) training a model via supervised fine-tuning (SFT) followed by another round of SFT that utilizes distillation on a chain-of-thought from a larger model Li et al. (2023); Hsieh et al. (2023). We confirm the effectiveness of the framework by fine-tuning two open models: Qwen2.5-3B Yang et al. (2024a) and Phi-4-Mini Microsoft et al. (2025), on the multiple choice question answering dataset MMLU-Pro Wang et al. (2024).
@@
For both steps of the pipeline, we provide sensitivity studies to the design choices made. In step one, we consider different options for complexity estimation based on a model’s confidence in its answer. Our detailed study considers a wide range of methods, including model-as-judge (MASJ), entropy-based aggregation, and calculation methods in line with findings from Fadeeva et al. (2023) with a top performing one being an answer entropy. In step two, we explore training on different complexity subsets of the distilled chain-of-thought — concluding that for regular data standard SFT is enough, while for hard data we benefit reasoning-based distillation. This observation constrains the reasoning to hard data only, making our approach token-efficient.
The specific contributions are the following:
@@
  • A training pipeline to fine-tune LLMs via SFT and distillation for regular and hard data correspondingly. Our procedure outperforms the standard SFT approach with accuracies $0.52$ and $0.64$ over two LLMs for MMLU-Pro c.t. $0.39$ and $0.51$ for SFT and $0.50$ and $0.63$ for distillation baselines.
  • An unsupervised method to split a multiple-choice question answering dataset by complexity based on the token-wise entropy of the response with ROC AUC $0.73$. Alternative complexity estimators rooted in MASJ and aggregated entropy, augmented with reasoning results and various aggregation techniques yield inferior results.
@@ 2 Related works
Complexity estimation.
Fadeeva et al. (2023) provides an in-depth comparison of multiple black-box and white-box methods of complexity estimation. They present promising results for white-box methods rooted in sequence probability and entropy aggregation. Sychev et al. (2025) focus specifically on entropy-based aggregations augmented with model-as-judge (MASJ) categorization by a reasoning score. They confirm that LLM’s token-level entropy of the output is a good predictor of question difficulty, especially in knowledge-based domains. They also introduce MASJ reasoning score to estimate the question complexity. However, the authors use these metrics only to analyze model behavior. They do not integrate it into a practical data aggregation or fine-tuning workflow.
Research gap.
The question if we can combine novel complexity estimation techniques based on entropy aggregates with adaptive learning methods remains open. It is also unclear how we can use the insights for the reasoning estimates to dynamically change our learning approach. This paper aims to address these gaps.
@@ 3.1 Training pipeline
We propose the complexity-aware fine-tuning pipeline detailed in Figure 1 with the following major stages: complexity estimation, data aggregation, fine-tuning.
@@
To aggregate the data into groups by complexity (regular, hard), we evenly divide the dataset into three parts, ordering the entries by entropy of the response. A group with lower entropy values is categorized as easy and with the higher values as hard, as Figure 2 depicts.
Figure 2: Data aggregation: we can further split the data to various complexity chunks
@@
For the regular groups, we use vanilla SFT Howard and Ruder (2018); Raffel et al. (2023), an established and robust practice. It involves fine-tuning a pretrained LM on labeled examples using a standard supervised objective, specifically like cross-entropy. Specific prompts are provided in Table 4 in Appendix.
As to the hard group, we hypothesize that as hard questions require multiple logical steps and multiple attempts to learn core facts. So, the standard SFT is suboptimal. We propose to elicit a chain-of-thought and allow the model to incrementally build the answer step-by-step as suggested by Wei et al. (2023).
In this work, we apply the distillation technique, which involves training a smaller student model on the chain-of-thought of a larger LLM. It is a well-known practice supported by Hsieh et al. (2023). To create the distillation training samples, we prompt a large LLM to answer the multiple-choice question and produce a chain-of-thought in the process. Next, the whole response is attached to the dataset and used to train the smaller model.
@@ 3.2 Complexity estimation approaches
To find the most suitable metric for the training pipeline, we analyze the performance of the following techniques: MASJ reasoning score, MASJ education level, a single token answer entropy, a chain-of-thought answer entropy, a chain-of-thought aggregated response entropy. Used prompts are available in Appendix A.1.
@@ 3.2.2 MASJ education level
3.2.3 Answer entropy
Single token answer entropy.
@@ 3.2.3 Answer entropy
where $p_{i}$ is the probability of a single token and $n$ is the vocabulary size. This is the main method used in our pipeline.
@@
Chain-of-thought answer entropy.
@@
3.2.4 Chain-of-thought aggregated response entropy.
Building upon the single-token entropy approach, we investigate more sophisticated methods in line with Fadeeva et al. (2023) for complexity estimation by analyzing the entire chain-of-thought (CoT) response. While the answer token entropy provides a localized measure of uncertainty, aggregating entropy across the complete reasoning process potentially offers a more comprehensive complexity assessment.
@@ 4.1 Experimental setup
We conduct all experiments on the multiple choice question answering dataset MMLU-Pro Wang et al. (2024), widely adopted by the community as one of the main performance benchmarks. It spans 14 domains, offering a broad selection of questions of varying complexities. Each question has approximately ten options, with a single correct one, which eliminates ambiguity in evaluation.
@@
Regarding the models  [^3] All models and datasets are published under permissive licenses that allow them to be used for research purposes, we utilize a variety of open models to analyze how the trend changes with model size. At the same time, we focus on smaller models for fine-tuning to make our results reproducible and more relevant to the distillation scenario.
We apply our pipeline to two models: Qwen2.5-3B and Phi-4-Mini. For them, we measure single token and chain-of-thought entropy, collect other response metadata, fine-tune the models, and evaluate the overall pipeline.
Bigger models are used to further extend the entropy aggregate and MASJ analysis: Mistral 24B, Phi-4, Mistral 123B Mistral (2024). For the chain-of-thought distillation we utilize an ensemble of models: DeepSeek-V3-0324 DeepSeek-AI (2024), Qwen 3 235B, Llama 4 Maverick Meta (2025).
4.1.1 Our method and baselines
In experiments, we performed training for 10 or 20 epochs, see more information below. In case of the reasoning used during training, we split training to two equal halves for the first and the second round of fine-tuning.
@@ 4.1.1 Our method and baselines
As our first baseline, we train the model via SFT without the data split for 10 epochs.
Curriculum
For the second baseline, we train the model via curriculum-based SFT: 3 epochs on easy data, 3 epochs on medium data, and 4 epochs on hard data.
@@
4.1.2 Technical details
Based on ROC AUC results and positive performance of SFT for medium and easy questions (see 4.3), we take single token entropy as our complexity metric for the pipeline described in Section 3.1.
4.2 Main results
(a) Accuracy for fine-tuning pipelines after 10 epochs for Qwen 3B
(a) Accuracy for fine-tuning pipelines after 20 epochs for Qwen 3B
(b) Accuracy for fine-tuning pipelines after 20 epochs for Phi-4-mini
Table  1 and Figures  3(a)  3(b) show the results. We see that the proposed training scheme results in a significant improvement over SFT, curriculum-based SFT and an alternative training scheme that uses distillation for only easy and medium questions. Qwen 3B achieves an accuracy of 0.50 compared to 0.47 and 0.42 for alternative and baseline, respectively, while Phi-4-mini gets to 0.60 compared to 0.58 and 0.46. At the same time, we see that the pipeline provides comparable performance to the distillation with significantly less data in number of tokens processed. Qwen 3B achieves an accuracy of $0.50$ compared to $0.49$ saving $79\%$ tokens, while Phi-4-mini gets to $0.60$ compared to $0.63$ saving $82\%$ tokens. We can also detect an uptrend in the pipeline training while the full distillation approach exhibits the signs of a plateau.
To confirm the plateau trend, we performed an extended set of experiments for 20 epochs with the same split of training epochs. Table 2 and Figures  4(a)  4(b) show the results. We find an even stronger confirmation of the previous findings with the proposed scheme outperforming all baselines. Qwen 3B achieves an accuracy of $0.52$ compared to $0.50$ and $0.39$ for distillation and baseline, respectively, while Phi-4-mini gets to $0.64$ compared to $0.63$ and $0.51$. The same data savings of $79\%$ and $82\%$ apply.
@@ 4.2 Main results
SFT 0.42 / 29k 0.55 / 27k
Curriculum 0.45 / 29k 0.54 / 27k
Distillation 0.49 / 19.72 k 0.63 / 15.15k
Alternative 0.47 / 5.88k 0.58 / 4.91k
Pipeline (ours) 0.50 / 3.99k 0.60 / 2.67k
@@
Method Qwen 3B Phi4-mini
SFT 0.39 / 59k 0.51 / 54k
Distillation 0.50 / 39.45k 0.63 / 30.30k
Pipeline (ours) 0.52 / 7.98k 0.64 / 5.35k
Table 2: Accuracy / tokens processed for complexity-aware fine-tuning pipelines after 20 epochs
4.3 Sensitivity study
4.3.1 Alternative complexity estimates
@@
The same data splitting procedure as described in Section 4.2 is applied for MASJ reasoning score and single-token entropy as complexity metrics.
@@
Figures 5(d) 5(c) show the results of SFT for each group split by the single-token entropy. For Phi-4-mini, we see that medium and easy questions outperform hard ones for the first 10 epochs. Shortly after, performance starts to decline for all groups. For Qwen 3B, we do not see a significant difference between the groups. Moreover, the performance plateaus after five epochs.
Figures 5(b) and 5(a) show the results of SFT for each group split by the MASJ reasoning score. We do not see a strong difference in performance between the groups. In combination with questionable ROC AUC scores, provided in Table 9, it makes MASJ reasoning score a less favorable metric for further experiments.
(a) Pipeline performance with cross-entropy (Qwen 3B)
(b) Pipeline performance with cross-entropy (Phi-4-mini)
Figure 6: Pipeline complexity metric performance comparison for entropy and cross-entropy across Phi-4-mini and Qwen 3B models.
Figures 6(a) and 6(b) show pipeline training results with cross-entropy as a complexity metric compared to entropy. We observe comparable results for Qwen 3B, while Phi4-mini demonstrates superior performance with entropy. The difference in performance could be attributed to entropy capturing the information across the entire distribution instead of a single correct token.
Figure 7: Curriculum learning accuracy dynamics for different models for Qwen 3B (left) and Phi-4-mini (right)
4.3.2 Curriculum learning fragility
Multi-stage pipelines such as curriculum learning show extreme sensitivity to hyperparameters such as a number of training epochs, learning rates and etc.
Figure 7 reveal the fragility of the original curriculum learning approach based on SFT. In this experiment we train a model on the data with increasing complexity for 10 and 20 epochs with different splits:
  • 10 epochs: 3 epochs on easy data, 3 epochs on medium data, and 4 epochs on hard data.
  • 20 epochs: 5 epochs on easy data, 5 epochs on medium data, and 10 epochs on hard data.
We see how Qwen 3B plateaus at a lower accuracy with more training epochs, while Phi4-mini quickly overfits on the data and its performance degrades.
4.3.3 Quality of complexity estimation
Method Qwen 3B Phi4-mini
Without split 0.72 / 0.70 0.72 / 0.74
Education level
High school and easier 0.73 / 0.72 0.76 / 0.75
Undergraduate 0.73 / 0.71 0.72 / 0.77
Graduate 0.66 / 0.65 0.64 / 0.68
Postgraduate 0.63 / 0.52 0.64 / 0.63
MASJ reasoning score
Low 0.72 / 0.71 0.78 / 0.79
Medium 0.72 / 0.70 0.70 / 0.72
High 0.64 / 0.62 0.59 / 0.58
(a) ROC AUC values for single-token response
Method Qwen 3B Phi4-mini
Answer Entropy (AE) 0.68 / 0.67 0.61 / 0.58
COT Mean 0.59 / 0.58 0.59 / 0.63
COT Max 0.63 / 0.61 0.6 / 0.65
Seq Mean 0.6 / 0.59 0.6 / 0.62
Seq Max-Mean 0.59 / 0.58 0.59 / 0.61
Seq Mean-Max 0.62 / 0.6 0.59 / 0.62
Marg Diff Mean 0.58 / 0.57 0.58 / 0.61
Top-2 Diff 0.51 / 0.5 0.5 / 0.51
COT Max - AE 0.54 / 0.53 0.51 / 0.57
COT Max + AE 0.7 / 0.69 0.62 / 0.62
(b) ROC AUC values for CoT response
Table 3: ROC AUC values for single-token and CoT responses
Second result is for the alternative prompt to allow model answer "I do not know"